3sqrt6(sqrt10-sqrt3)
Answer by Boreal(15235)

You distribute the 3 sqrt (6), and you recognize that sqrt (6) * sqrt (10)= sqrt (60)
The problem is 3 sqrt (60)- 3 sqrt (18)
Then you try to factor a perfect square out. sqrt (60)=sqrt (4)*sqrt(15)=2 sqrt (15)
sqrt (18)= 3 sqrt (2)
So the second line become 6 sqrt (15)-9 sqrt (2). These cannot be simplified, although one could factor out a 3[2 sqrt (15)-3 sqrt (2)]
Calculator 7.348* (3.16-1.73)=23.22-12.71
6 sqrt (15)-9 sqrt (2)=23.24-12.73 ; These are basically the same with rounding error.